Find the positions at which a particle of mass m move on the x-axis with energy v=E/q^4(x^4+4ax^3-8a^2x^2) is in stable equilibrium
1
Expert's answer
2020-01-22T04:23:50-0500

"V(x)=\frac{E}{a^4}(x^4+4ax^3-8a^2x^2)"

*q isn’t given, so I think there should be a (in this case we know, that a ≠ 0)


For the equilibrium positions, "\frac{d V}{d x }" will be zero.

"\frac{d V}{d x}=\frac{d}{d x}(\frac{E}{q^4}(x^4+4ax^3-8a^2x^2))=\frac{E}{q^4}(4x^3+12ax^2-16a^2x)=0"

"\frac{E}{a^4} 4x(x^2+3ax-4a^2)=0"

"x(x+4a)(x-a)=0"

"x_1=0, x_2=-4a, x_3=a" (if a≠0, then all these points are different)


For stable equilibrium "\frac{d ^2 V}{d x^2}" must be greater than zero so that potential energy is a minimum.

"\frac {d^2 V}{d x ^2}=\frac{d}{dx}(\frac{E}{a^4}(4x^3+12ax^2-16a^2x))=\frac{E}{a^4}(12x^2+24ax-16a^2)"

At "x=0" gives "\frac{d ^2 V}{d x^2}=-\frac{16E}{a^2}"

At "x=-4a:" "\frac{d ^2 V}{d x^2}=\frac{80E}{a^2}"

At "x=a: \frac{d ^2 V}{d x^2}=\frac{20E}{a^2}"


Answer:

if "E>0: \; x_1=-4a,x_2=a" are positions of stable equilibrium

if "E<0: \;x=0" is the point of stable equilibrium

if "E=0 \Rightarrow V(x)=0" neutral equilibrium.